OpenStudy (anonymous):

1. Renting a bicycle at a local Bike shop costs $30 for the first day and $10 for each additional day. a. Create a chart of values and write an equation to find the total cost, C in terms of the number of days d. b. Use the equation you created in Part a to find the cost of renting a bike for 5 days. 2. Simplify the following problem using the distributive property: 2x + 3y + 2(x – y) – 3x. Show each step in your calculation. Justify each of your steps using number properties. Answer: Please help will become a fan and give you a medal :)

OpenStudy (anonymous):

@phi

OpenStudy (phi):

which question ?

OpenStudy (anonymous):

both please

OpenStudy (dragonbreath):

1.c=10d+30 d is days and c is cost c=80

OpenStudy (dragonbreath):

2x+3y+2(x-y)-3x= =2x+3y+2x-2y-3x= =x+y